OpenStudy (isabelle99):

A geometric sequence has a4=4 and a5=7 . What is a1 ?

OpenStudy (isabelle99):

please help... I want to say a1= -5 but i think i did the arithmetic sequence way and not the geometric sequence way

OpenStudy (salty):

Do you know how to write the general term of a GP

OpenStudy (isabelle99):

no

OpenStudy (salty):

Its written like this \[T_n=ar^{n-1}\] \(T_n\) =\(n^{th} \) term a= 1st term r=common ratio

OpenStudy (isabelle99):

But i don't have the common ratio nor do i have a1

OpenStudy (salty):

You can get the common ratio by dividing two consecutive terms

OpenStudy (salty):

For example Common ratio=(2nd term) /(1st term)

OpenStudy (isabelle99):

but i only have the 4th term and 5th term do i use those??

OpenStudy (salty):

The 4th and 5th terms are consecutive You get the 5th term when you multiply the 4th with common ratio Dividing the 5th term by the 4th one will give you the common ratio

OpenStudy (isabelle99):

so 1.75 would be r???

OpenStudy (salty):

Yes

OpenStudy (salty):

Next we need to find "a"

OpenStudy (isabelle99):

would a3 be 2.29 or did i do that wrong??

OpenStudy (salty):

Correct! :)

OpenStudy (salty):

See that equation i wrote above When you plug \(T_n\) =4, n=4 and r=1.75 you can find a

OpenStudy (salty):

a is our 1st term tho and thats what we need to find

OpenStudy (salty):

btw you can also find a( or you can call it a1) by same way you found a3

OpenStudy (isabelle99):

i got a1=0.75 but just just kept dividing the number i got by 1.75 until i got to a1...i didn't use the formula i still don't understand it

OpenStudy (salty):

Thats correct but your approach might not be suitable if the question asked to find a53 cuz then it will be too much calculation I can help u understand the fornula

OpenStudy (isabelle99):

ok please explain

OpenStudy (salty):

Suppose the 1st term is a and common ratio is r You get the a2 by multiplying a with r right So a2 will be - > ar Similarly you can get ae by multiplying r and a2 So a3 =ar^2 a4=ar^3 a5=ar^4 a6=ar^5 a7=ar^6 a8=ar^7 See a pattern forming here? The \(n^{th} \) term is given by \(T_n = ar^{n - 1}\)

OpenStudy (salty):

That not ae Its supposed to be *a3

OpenStudy (salty):

a4=ar^3 = ar^(4-1) a5=ar^4 =ar^(5-1) a6=ar^5 =ar^(6-1) a7=ar^6 =ar^(7-1) a8=ar^7 =ar^(8-1)

OpenStudy (isabelle99):

okay but this one started at a4 wouldn't i have to work backwards to get to a1

OpenStudy (salty):

You found r and you already know the value of a4 And you also know that \[a_n=ar^(n-1)\] Btw \[T_n\] and \[a_n\] are the same thing For a4 - n will be 4 r is 1.75 and a4 is 4 Soo \[4=a(1.75)^{4-1}\] solve this to get a

OpenStudy (isabelle99):

i understand it thank you and what if i wanted a2 instead of a1? what would i had plug in differently?? (im not asking you to work the whole thing out just plug it in for me)